OpenStudy (anonymous):

I need help, i'm confused? I will be giving best answer to the person who I find the most helpful. Help me find the value of the variable that makes the following statements true? I will be giving medal to best answers. <3 Fifteen is equal to ten less than a number. The sum of a number and six is equal to the sum of two times the number and one.

OpenStudy (anonymous):

Would the first be N-10=15?

OpenStudy (anonymous):

Would the second be? N+6=N*2+1

OpenStudy (anonymous):

If not, please tell me where I am wrong?

OpenStudy (campbell_st):

you are correct in both the 2nd equation can be written as N + 6 = 2N + 1

OpenStudy (anonymous):

Thankyou so much, haha. Now I know how to do verbal phrase equations yay.